Solve the equation by factoring
6x(x − 1) = 7 − 25x

Answer by mananth(16946)

6x^2-6x=7-25x
6x^2+19x-7=0
6x^2+21x-2x-7=0
3x(2x+7)-1(2x+7)=0
(2x+7)(3x-1)=0
2x+7 =0
x= -7/2
OR
3x-1=0
x=1/3